OpenStudy (anonymous):

7/ radical 3x = radical (x-4)

OpenStudy (anonymous):

\[7\div \sqrt{3x} + \sqrt{x-4}\]

OpenStudy (anonymous):

It's a PLUS!! Sorry!

OpenStudy (phi):

\[\frac{7}{\sqrt{3x}}+ \sqrt{x-4}\]

OpenStudy (anonymous):

This is what I got when I multiplied by the conjugate - how close am I? \[7\sqrt{3x}- 7\sqrt{x-4}\div2x-4\]

OpenStudy (phi):

I suppose they want you to get rid of the square root in the denominator

OpenStudy (anonymous):

yes that is correct

OpenStudy (phi):

I don't think we want to do a conjugate

hero (hero):

Um....you need someone to go over this with you in detail.

OpenStudy (anonymous):

I understand that, but I have no one until after the assignment is due - zero period 7am tomorrow! Any help is appreciated!

hero (hero):

I'd love to help you out but I charge a fee for personal tutoring

OpenStudy (phi):

When you have a fraction, and the denominator is a square root, multiply top and bottom by the square root.

OpenStudy (phi):

if you have a fraction, and the bottom is a square root PLUS a number, multiply the top and bottom by the conjugate.

OpenStudy (anonymous):

Thanks, but I don't have any money, and my parents are on a single income. I will do my best to understand what information is hared wuith me at this time.

OpenStudy (anonymous):

@phi That's what I did, but I'm not sure the mechanics are all there

hero (hero):

How many questions do you have?

OpenStudy (anonymous):

That's how I ended up with the radicals in the numerator. basically a 7 infromnt of each portion of the equations conjugate.

OpenStudy (anonymous):

@ hero This is my last one!

OpenStudy (phi):

Just to make sure I understand the problem is it \[\frac{7}{\sqrt{3x}}+ \sqrt{x-4}\]

OpenStudy (phi):

or is it \[\frac{7+ \sqrt{x-4}}{\sqrt{3x}}\]

OpenStudy (anonymous):

@ phi no the x-4 under the radical is attached to the radical 3x - they are BOTH below the division bar

OpenStudy (phi):

\[\frac{7}{\sqrt{3x}+ \sqrt{x-4}}\]

OpenStudy (anonymous):

YES!!!

OpenStudy (anonymous):

How do you do that? I can not figure this thing out!

OpenStudy (phi):

It's tough answering the question when I'm doing the wrong one!

hero (hero):

lol, wow! Phi great job clarifying

OpenStudy (phi):

so you do use a conjugate

OpenStudy (phi):

so what did you get for the bottom?

hero (hero):

@ jenn: Next time, write 7 OVER (sqrt{3x} + sqrt{x - 4})

OpenStudy (anonymous):

i got 2x-4 fpr the denominator, and \[7\sqrt{3} - 7\sqrt{x-4}\] for the numerator

OpenStudy (anonymous):

@hero Thanks!

OpenStudy (anonymous):

am I close?

OpenStudy (phi):

should be 3x - (x -4)= 3x-x +4 = 2x+4

OpenStudy (phi):

top looks good assuming there is an x in the first square root with the 3

OpenStudy (anonymous):

for the denominator? - Yes I accidentally forgot the x on the computer, but it's on my paper!

OpenStudy (phi):

numerator

OpenStudy (anonymous):

Yes - Thanks! I have log questions next! Any interest in helping? I will post them on the other side!

OpenStudy (phi):

Until you stump me

OpenStudy (anonymous):

You are so awesome! I posted it already - give it a look and let me know where to start!